### v3

# VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

Abstract
Most existing benchmarking approaches for evaluating the output quality of large language models (LLMs) rely on comparing LLM responses to predefined references. Such methods, based on static datasets, quickly become outdated as LLM capabilities and use cases evolve. In this work, we introduce VARCO Arena—a novel, cost-effective, and robust benchmarking approach that leverages a single-elimination tournament structure to minimize the number of required comparisons while eliminating the need for static references or costly human annotations. We validate our approach through two experiments: (i) a simulation study that examines its robustness under various conditions, and (ii) an empirical evaluation using publicly available benchmark prompts. In both experiments, VARCO Arena consistently outperforms current LLM benchmarking practices, achieving stronger correlations with human-established Elo ratings. Our results demonstrate that VARCO Arena not only produces reliable LLM rankings but also provides a scalable, adaptable solution for qualitative evaluation across diverse, customized use cases. We release our demo and code at [https://huggingface.co/spaces/NCSOFT/VARCO_Arena].
VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models
Seonil Son1, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: deftson@ncsoft.com
1 Introduction
The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks. The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios Ouyang et al. (2022); Köpf et al. (2024); Roziere et al. (2023). As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant. In response to this challenge, numerous benchmarks evaluating LLM capabilities have emerged Hendrycks et al. (2020); Srivastava et al. (2023); Zhong et al. (2024). Many LLM benchmarks employ formats amenable to automated scoring. Examples include benchmarks for arithmetic problems Gao et al. (2022); Patel et al. (2021), multiple-choice questions Lin et al. (2022), and code execution Austin et al. (2021); Chen et al. (2021). While these benchmarks are valuable, they primarily assess problem-solving abilities. The need for LLM benchmarks extends beyond this, as the primary value of LLMs lies in the versatility of their generative behavior. Researchers have primarily relied on pairwise comparisons to evaluate and rank the LLMs based on human preference annotations. A representative example of this approach is Chatbot Arena Chiang et al. (2024), which computes Elo ratings based on a massive number of human votes. While Chatbot Arena benefited from reliable, open-ended prompts generated by a large user base, obtaining such a high volume of annotations from dedicated annotators remains challenging.
Figure 1: VARCO Arena directly compares LLM response pairs in single-elimination tournament rather than comparing references. In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
No. Comp. ($\downarrow$) Judge Eval. Type
Chatbot Arena unknown human head-to-head
Current Practice $n_{\text{model}}\cdot|X|$ LLM reference -based
VARCO Arena $(n_{\text{model}}-1)\cdot|X|$ LLM head-to-head
Table 1: Comparison between Current Practice and Varco Arena. $|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively. Human annotators are considered much more costly than LLM judge counterpart.
To address these challenges, several benchmark datasets for generation tasks have been developed and applied with an LLM judge Zheng et al. (2024) to quantify LLM capabilities. Using these benchmarks and their reference responses, LLMs are ranked via pairwise comparisons between their outputs and the reference responses provided by LLM judges. Representative examples include AlpacaEval Li et al. (2023), Arena-Hard-Auto Li et al. (2024), and MTBench Zheng et al. (2023). There are two major advantages to relying on reference responses: (1) The number of comparisons required for ranking scales linearly with the number of LLMs. (2) The reference response establishes a quality standard for evaluating LLM outputs.
However, we argue that using reference responses to mediate comparisons between LLM outputs is suboptimal and that head-to-head comparisons yield more accurate assessments. In this context, we propose VARCO Arena as a novel evaluation framework that directly compares LLM responses while requiring fewer comparisons. Even with fewer comparisons and without relying on reference responses, VARCO Arena achieves a stronger correlation with the human-established Elo rankings from Chatbot Arena.
VARCO Arena is designed to conduct a single-elimination tournament for each prompt across all participating LLM responses and then compute Elo ratings Elo and Sloan (1978) based on the outcomes of these matches. By organizing tournament-style matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs with fewer comparisons than those required by reference-based methods. These win rates are subsequently interpreted as Elo ratings, enabling us to generate a comprehensive ranking of candidate LLMs.
We validate the effectiveness of VARCO Arena against current LLM benchmarking practices through two experiments. First, we conduct a simulation study (Section 4.2) to evaluate the reliability and robustness of our tournament approach under various controlled factors, including the number of participating LLMs, the number of test prompts, and the accuracy of the judge model. Second, we empirically demonstrate that running VARCO Arena with several judge LLMs on public benchmark (Section 4.3) prompts yields a stronger correlation with human-established Elo rankings from Chatbot Arena than does the reference-based approach.
These experiments confirm that our benchmarking approach outperforms current practices in terms of ranking reliability while requiring the lesser number of comparisons (Table 1). Notably, the simulation results—which emulate outcomes under various conditions—underscore that the superiority of our tournament approach is not coincidental but rather reflects the inherent advantages of direct pairwise comparisons over mediated comparisons using reference responses.
In this work, we make the following contributions:
  1. 1.
    We introduce VARCO Arena, a novel reference-free benchmarking method for LLMs that leverages a single-elimination tournament structure to perform head-to-head comparisons of LLM responses across test prompts. This approach eliminates the dependency on static references and enhances benchmarking flexibility.
  2. 2.
    Through extensive simulation studies and empirical evaluations, we demonstrate that VARCO Arena not only yields more reliable LLM rankings compared to traditional reference-based methods but also achieves this with significantly fewer comparisons, thereby reducing evaluation costs.
  3. 3.
    We release our demo and code to facilitate the further research and reproducibility; our resources are available at [https://huggingface.co/spaces/NCSOFT/VARCO_Arena]
2 Preliminaries: Quantifying Generation Ability
Quantifying an LLM’s generation ability is crucial, but it presents several challenges. The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as the provided test prompts and the inherently subjective nature of human preferences. A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its real-world performance. Two popular measures for LLM generation ability are the win rate against reference responses and the Elo rating.
2.1 Win Rate in Comparison to the Reference
AlpacaEval Li et al. (2023) and Arena-Hard-Auto Li et al. (2024) are representative benchmark datasets that aim to quantify LLM response quality using reference responses. These benchmarks employ an LLM judge with a standardized prompt for automated evaluation of generation capability. Given the prompts, LLM responses, and reference responses, the judge LLM is tasked with determining whether the reference or the response is preferred. The LLM’s win rate across the test prompts in the benchmark is used as a measure of its generation proficiency.
2.2 Elo Rating
The Elo rating system, introduced by Elo and Sloan, has since become a popular method for quantifying performance levels in competitive sports and, more recently, for evaluating the generative abilities of LLMs. The primary purpose of the Elo rating system is to represent a participant’s skill level as a single scalar value, enabling the prediction of relative win rates between participants who have never directly competed. Assessing the superiority of one LLM over another in terms of generative capability shares several similarities with determining winners in competitive sports. Varying test prompts may yield different results, similar to how weather or other factors can influence the outcome of a sports match. While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently. With Elo ratings, one can expect the relative chance of winning between a pair of players. The expected win rate of the player $i$ against $j$ is computed as follows:
[…]
Computing accurate Elo ratings requires a sufficient number of matches to estimate relative win rates between pairs of participants. The resulting matrix of relative win rates is then used to estimate Elo ratings through logistic regression, similar to multivariate logistic regression with a sigmoid function. In Equation 1, $P(i>j)$ represents the expected win rate of participant $i$ over $j$, $R_{i}$ is the Elo rating of participant $i$, and $\Delta_{\text{ij}}$ is the Elo rating difference between participants $j$ and $i$ ($R_{j}-R_{i}$).
Chatbot Arena presents a leaderboard of LLMs, with Elo ratings computed from matches evaluated by a large user base. Users prompt a pair of LLMs and submit their judgments of which one responded better. Although Chatbot Arena relies on manual evaluation of matches, it offers an intuitive method for comparing LLMs using Elo ratings.
3 VARCO Arena
We propose VARCO Arena, a reference-free approach for benchmarking large language models (LLMs) via single-elimination tournaments. Instead of comparing model responses against fixed references, VARCO Arena directly compares outputs from different models, determining superiority through head-to-head matchups for each prompt in our benchmark dataset. Repeated tournaments across prompts yield reliable leaderboards that reflect the relative performance of each model.
We begin by motivating our approach over current reference-based evaluations (Section 3.1). Next, we detail how VARCO Arena performs tournaments and results in Elo ratings (Section 3.2 and Algorithm 1). Finally, we discuss the reason why aggregating the results from tournaments promises reliable ranking (Section 3.3).
3.1 Comparing to a Reference is not Always Helpful
Although reference texts are a standard way to evaluate and rank large language models (LLMs), they introduce potential failure modes. Beyond the fact that a single reference might not capture every dimension of correctness, relying solely on a reference can lead to unreliable rankings of LLMs.
Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs. If we choose to compare LLM responses directly to rank them using Elo ratings (Equation 1), all head-to-head comparisons are utilized. In contrast, reference-based evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ \quad\text{vs.}\,\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\hfil\hskip 20.00003pt&\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})>Y_{i}&\text{(unhelpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})<Y_{i}&\text{(unhelpful)}\end{cases}\end{array}$
When the reference output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the reference is effective for benchmarking. Otherwise, these comparisons do not help differentiate LLM performance. Consequently, the reference-based approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the reference.
3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
Algorithm 1 Tournaments of LLMs over prompts
1:prompts $X=\{x_{1},~{}x_{2},~{}...,~{}x_{i}\}$, LLMs $M=\{m_{1},~{}m_{2},...,~{}m_{j}\}$, outputs $O_{i,j}$ = $m_{j}(x_{i})$
2:Ranked LLMs with Elo ratings
3:function Match($m_{1},m_{2},x$)
4:     return $m_{1}$ if $\texttt{IsBetter}(O_{x,1},O_{x,2})$
5:     else $m_{2}$
6:end function
7:function SingleElim($M$, $x$, res)
8:     if $|M|=2$ then
9:         res.append(Match($M$[0], $M$[1], $x$))
10:         return res[-1]
11:     end if
12:     mid $\leftarrow\lfloor|M|/2\rfloor$
13:     left $\leftarrow$ SingleElim($M$[:mid], $x$, res)
14:     right $\leftarrow$ SingleElim($M$[mid:], $x$, res)
15:     return SingleElim(left + right, $x$, res)
16:end function
17:function Tournaments2Ranks($X,M$)
[…]
19:     for $x_{i}\in X$ do
20:         SingleElim(Shuffled($M$), $x_{i}$, res)
21:     end for
22:     return ComputeElo(res)
23:end function
Figure 1 and Algorithm 1 illustrate how VARCO Arena benchmarks LLMs via a tournament approach. Here, $|X|$ denotes the number of prompts in the benchmark dataset. Each execution of VARCO Arena runs a tournament among participant LLMs for every prompt in the dataset.
The use of tournament structures for LLM benchmarking offers both benefits and challenges. A major advantage of a single-elimination tournament is efficiency. As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using references. However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
To retain tournament’s efficiency while obtaining a fine-grained ranking, we propose aggregating tournament results over multiple prompts with randomized initial match-ups for each prompt. Performing multiple tournaments with random initialization offers several benefits:
  1. 1.
    It resolves ties among non-champion participants from previous tournaments.
  2. 2.
    It mitigates the impact of unfavorable match-ups in any single tournament.
  3. 3.
    Aggregating match results allows for precise win rate estimation via Elo ratings, resulting in a well-aligned overall ranking.
  4. 4.
    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt.
In Section 3.3, we further explain how aggregating multiple tournaments could yield an reliable ranking of LLMs. We also provide an analysis of the number of matches each LLM faces, offering a comprehensive view of the method’s efficiency and effectiveness.
3.3 Why Aggregating Multiple Tournaments Yields Reliable Ranks?
Our approach aggregates match outcomes from multiple tournaments to approximate the complete set of pairwise comparisons—akin to the comparisons made in merge sort. In a single-elimination tournament, every participant advances based solely on match outcomes, a process that mirrors the merging steps in merge sort. Notably, a single-elimination tournament executes only the comparisons strictly necessary for determining the winner, omitting many comparisons that merge sort would perform further.
We posit that the missing pairwise match-ups in any one tournament can be recovered by aggregating tournaments conducted over different prompts. This hypothesis relies on the assumption—central to our use of the Elo model—that match outcomes are independent of the prompt. Consequently, matches across different prompts are considered equivalent.
Considering only the initial comparisons, which are randomly sampled, the aggregate number of comparisons is $|X|\cdot n_{\text{model}}/2$. Since $n_{\text{model}}$ is typically on the order of tens and $|X|$ comprises at least hundreds, this number exceeds the total possible match-ups, $\binom{n_{\text{model}}}{2}$ [^1] Note that merge sort produces a set of match-ups with no duplicates.. Furthermore, as shown in Table 1, the remaining matches—totaling $|X|\cdot(n_{\text{model}}-1)$—either contribute additional merge sort structures or enhance the accuracy of estimating relative win rates among LLM participants.
Moreover, considering that each unique pair meets in at least $|X|/(2(n_{\text{model}}-1))$ matches [^2] A model participates in between $|X|$ and $\lceil\log_{2}n_{\text{model}}\rceil$ matches per tournament. Dividing this by the number of possible match-ups per model, $2(n_{\text{model}}-1)$, yields the expression above. across the benchmark, this frequency is sufficient for a fair estimation of their relative win rates.
In summary, aggregating tournaments not only reconstructs the full set of pairwise comparisons for a merge sort but also ensures that each pair of models faces one another often enough to yield accurate win rate estimations, leading to reliable Elo ratings. Based on these considerations, we propose that conducting tournaments over a benchmark prompts will yield a reliable ranking of LLMs.
4 Experiments
We run two experiments with different settings for comparing VARCO Arena and current practice of reference-based benchmarking. In the first experiment (Section 4.2) we conduct a simulation study to control various factors that affect LLM benchmarking. This simulation tests our foundational propositions for VARCO Arena design (mentioned in Section 3.1 and 3.3) under a more controlled, simplified environment immune to noisy factors such as potential biases of LLM judges Park et al. (2024).
The other is empirical experiments (Section 4.3). We use gpt-4o and -mini as well as other several popular models such as Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to validate the effectiveness of our tournament approach against current LLM benchmarking practices. By presenting both simulation and empirical results, we aim to demonstrate the effectiveness of our tournament approach. In Section 4.1, we first describe the common experimental settings before getting into specific details of each experiments in the following subsections.
4.1 Chatbot Arena Leaderboard Ratings as Ground-Truth LLM Rankings
We compare the results of each benchmarking approach against the rankings from the Chatbot Arena leaderboard. Chatbot Arena is widely regarded as one of the most reliable leaderboards due to its extensive collection of human preference annotations. Given the large number of votes and the diverse set of prompts used for model comparisons, the resulting rankings are considered sufficiently accurate to serve as ground truth.
4.2 Experiment 1: Simulation Study
We designed a simple simulation to emulate a probabilistic model of LLM matches, adhering to the Elo rating system in a controlled environment. In line with the Elo model’s assumptions, our judge is configured to follow Equation 3 exactly. The judge stochastically determines the winner of each LLM match solely based on the Elo rating difference ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$). As described in Equation 3, the outcome of a single match is sampled according to $P_{\text{predict}}(i>j)$, which is computed as the product of the judge’s accuracy and the likelihood of model $i$ beating model $j$ based on the Elo gap ($P_{gt}$).
$\begin{split}P_{\text{predict}}(i>j)&=P_{\text{judge}}\times P_{\text{gt}}(i>j)\\ &=P_{\text{judge}}\times\frac{1}{1+10^{\Delta_{\text{ij}}/400}}\end{split}$ (3)
The details of our simulation settings are as follows:
Ground-truth Elo ratings (initial parameter of the simulation, and at the same time, the gold ranking to reproduce): We extracted Elo ratings from the English category of Chatbot Arena as of June 23. This Elo ratings are the estimates from massive user-submitted judgments (approximately 60% of the total submissions to the platform). For the simulation, any set of Elo ratings could be used, but we opted for real values computed from human preferences.
Judge Accuracy ($P_{\text{judge}}$): In practice, a judge’s accuracy is an adaptive value that depends on factors such as the specific prompt-response pair and the manner in which LLM judges are prompted. In this simulation study, we control this parameter, varying it from 0.6 to 0.9.
Number of Participant LLMs ($n_{\text{model}}$) and Benchmark Dataset Size ($|X|$): We varied these factors to assess the robustness of both the tournament and reference-based approaches under diverse conditions. This allowed us to explore how reliability changes with the number of participants in both data-poor and data-rich environments.
Simulation Procedure:
  1. 1.
    Select the participant LLMs and obtain their Elo ratings for the simulation.
  2. 2.
    Compute the expected relative win rates ($P_{\text{gt}}$, see Equation 3) from the participants’ Elo ratings.
  3. 3.
    Sample match outcomes for each LLM pair according to the benchmarking approach. The winner of each match is sampled from $P_{\text{pred}}$ (Equation 3), which depends solely on the Elo gap ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$).
  4. 4.
    Repeat step 3 for the designated number of test prompts ($|X|$).
  5. 5.
    Calculate scores for ranking:
    1. (a)
      For the reference-based approach (current practice), use the win rate against the reference model (gpt-4-1106-preview, which has an Elo rating of 1233).
    2. (b)
      For the tournament approach (VARCO Arena), compute Elo ratings from all match outcomes.
  6. 6.
    Rank the models based on these scores.
  7. 7.
    Compute the Spearman correlation between the simulated rankings (from step 6) and the ground-truth rankings (from step 1).
We perform 50 trials for each simulation configuration to mitigate randomness from tournament brackets and sampling.
4.3 Experiment 2: Assessing VARCO Arena Using Various LLM Judges
To empirically validate our proposal, we evaluated the reliability of both VARCO Arena and reference-based approach over the top 20 models from the Chatbot Arena leaderboards. This experiment employs actual prompt inputs and LLM outputs, distinguishing it from the earlier simulation study.
4.3.1 Dataset: Test Prompts and LLM Responses Used
Testing the benchmarking approaches requires: (1) test prompts and (2) the corresponding responses from LLMs. For the benchmark dataset, we selected Arena-Hard-Auto Li et al. (2024). The prompts in Arena-Hard-Auto were carefully curated from Chatbot Arena user queries. This dataset consists of 500 prompts—two instances for each of 250 subtopics. Although AlpacaEval Li et al. (2023), which comprises 800 prompt-reference pairs, could serve as a viable testbed, we opted for Arena-Hard-Auto because its design aligns more closely with Chatbot Arena. Arena-Hard-Auto uses responses from gpt-4-0314 as the reference outputs. For ranking, we utilized the reserved outputs of the top 21 models from the Arena-Hard-Auto Browser. [^3] Extracted from the 2024 Jul 6 commit (fd42026).
4.3.2 Participant LLMs
For ranking, we selected 20 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
4.3.3 LLM Judges
We used several aligned LLMs as judges for testing both benchmarking approaches. LLMs of our choice are gpt-4o family of models OpenAI et al. (2024), Claude3.5, and a selection of open-weight models: Qwen2.5 Qwen et al. (2025), Llama3.1 Grattafiori et al. (2024), and Gemma2 Team et al. (2024). For pairwise comparisons of responses, we employed the judging prompt suggested in LLMBar Zeng et al. (2024) (See Appendix A.6.2). The same judge prompt was applied consistently across both the tournament and reference-based approaches. To mitigate position bias Wu and Aji (2023), the order of model responses was alternated during evaluation. Further details on the LLM-as-a-judge configuration are provided in Appendix A.6.
The two experimental settings are summarized as follows:
Experiment 1 (Simulation Study): This experiment uses the ground truth Elo ratings of the models to initialize the simulation. We vary control parameters for the benchmarking approaches—including the judge’s accuracy ($P_{\text{judge}}$), the number of test prompts used ($|X|$), and the number of participant LLMs ($n_{\text{model}}$)—to determine which benchmarking approach more accurately reproduces the participants’ ranking. For each configuration, we conduct 50 trials of experiments.
Experiment 2 (Empirical Runs): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges. We select the top 20 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts. For both the tournament and reference-based approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings. In our empirical study, we conduct 500 trials for each experimental setting.
Figure 2: Simulation results comparing the tournament and reference-based approaches. The tournament method consistently outperforms the reference-based approach in Spearman correlation across various control variables: the number of participant LLMs ($n_{\text{models}}$), the number of benchmark prompts ($|X|$), and judge precision ($P_{\text{judge}}$).
5 Results and Discussion
We assess the reliability and robustness of VARCO Arena as a means for LLM benchmarking, comparing it against the current reference-based approach. Our results from both simulation study and empirical runs indicate that the tournament approach of VARCO Arena yields rankings that align more closely with the ground-truth Elo leaderboards. We present our findings using whisker plots and tables in the following sections.
5.1 Experiment 1: Simulation Study Results
Figure 2 illustrates noticeable differences in Spearman correlation, indicating that the tournament approach is more reliable than the reference-based method. The consistent performance gap across various conditions—namely, the number of participants, the number of test prompts, and judge accuracy ($n_{\text{model}}$, $|X|$, and $P_{\text{judge}}$)—demonstrates the robustness of the tournament approach. Although the simulation simplifies real-world complexity, a similar performance gap was observed in the empirical findings (Experiment 2, Figure 3). This consistency suggests that the robust performance of VARCO Arena is not coincidental or limited to a specific empirical setting of ours.
5.2 Experiment 2: Empirical Results
Figure 3: Results of VARCO Arena (tournament) and reference-based approach with gpt-4o (left) and gpt-4o-mini (right) judge. VARCO Arena constantly records higher Spearman correlation coherent with the Experiment 1 result (Figure 2). Results summary is on Table 2.
As hinted in the previous section, the empirical results in Figure 3 show that VARCO Arena consistently outperforms the reference-based approach. Although the performance gaps are less pronounced than in the simulation, the same trend persists. In Table 2, we report the median values for VARCO Arena and the reference-based approach using the gpt-4o family of judges while varying the number of test prompts ($|X|$). These results consistently demonstrate that VARCO Arena outperforms the reference-based method. Note that VARCO Arena shows similar or superior reliability even in extreme data-poor benchmark condition ($|X|=50$).
Table 3 presents the outcomes when using other LLMs as judges, with a fixed number of prompts ($|X|=500$). The results for Claude3.5-sonnet, Llama3.1-8b, and Qwen2.5-7b follow a similar trend. However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable for benchmarking. Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using VARCO Arena or considering reference-based approach.
Spearman corr. ($\uparrow$) $|X|$ = 50 100 250 475 500
comp. to ref. (4o) 0.895 0.935 0.963 0.966 0.964
tournament (4o) 0.905 0.940 0.960 0.970 0.970
comp. to ref. (4o-mini) 0.895 0.908 0.917 0.916 0.912
tournament (4o-mini) 0.901 0.919 0.931 0.933 0.933
Table 2: Spearman correlation ($\uparrow$) varying over size of the benchmark set ($|X|$) for each benchmarking approach. Comp. to ref. refers to reference-based approach.
$|X|$ = 500 claude3.5 sonnet llama3.1 8b-it qwen2.5 7b-it qwen2.5 0.5b-it gemma2 2b-it
comp. to ref. 0.924 0.820 0.756 0.089 0.592
tournament 0.930 0.850 0.811 -0.124 0.552
Table 3: Spearman correlation ($\uparrow$) result using other LLMs as a judge. Comp. to ref. refers to reference-based approach.
5.3 Incorporating a New LLM into an Existing Leaderboard
While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case for leaderboards. We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a reference. Our findings indicate that the latter approach is more reliable (Table 4). Further details and discussions are provided in Appendix A.4.
$\bar{|\Delta_{\text{rank}}|}$ ($\downarrow$) gt=1-6 7-13 14-19 (20) total avg.
binary search (4o) 0.92 1.84 2.13 1.72
comp. to 1st (4o) 1.98 1.55 1.57 1.39
binary search (4o-mini) 1.27 1.82 1.21 1.5
comp. to 1st (4o-mini) 1.00 1.43 1.43 1.37
Table 4: Comparison of the binary search method versus using the top-performing model’s response as a reference (comp. to 1st) for inserting a new LLM into the leaderboard. We report the mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from the ground-truth leaderboard as an additional error metric. For further details, see Algorithm 2 in Appendix.
6 Related Works
6.1 Elo-based LLM Benchmarking
Recent studies have leveraged Elo ratings derived from human preferences to benchmark LLMs Boubdir et al. (2023). For instance, Chatbot Arena Chiang et al. (2024) employs Elo as an evaluation metric, while RAGElo Rackauckas et al. (2024) uses it as a ranking metric. These implementations highlight the benefits of applying Elo ratings for open-ended text quality evaluation.
6.2 Reference-free Evaluation
Reference-free evaluation has emerged as a promising alternative to static reference-based methods from the era of Natural Language Generation. Advances in LLM capabilities have enabled models to assess open-ended responses effectively Jauhiainen and Guerra (2024). When reference quality is poor, reference-free metrics like XComet Guerreiro et al. (2023) works as a more reliable alternatives.
7 Conclusion
We introduced VARCO Arena, a reference-free LLM benchmarking approach that employs a tournament-style framework with direct pairwise comparisons to evaluate the generative capabilities of LLMs. VARCO Arena offers a cost-effective, scalable, and adaptable solution for benchmarking response quality. Our simulation study and empirical experiments demonstrate that VARCO Arena consistently achieves higher rank reliability compared to current reference-based methods, as evidenced by stronger correlations with human-established Elo ratings. Given its robust performance and flexibility, we believe VARCO Arena can serve as a reliable automated tool for model selection and ranking across diverse and evolving use cases. Future work will explore broader applications, such as benchmarking LLMs in multi-modal settings (e.g., those incorporating visual or audio inputs and outputs).
Limitations
Although we tested robustness of the tournaments performed by VARCO Arena, it still has added factor for randomness such as match bracket initialization which does not applies to reference-based method. Rooms for improvement exist for more informative match-making algorithm that would achieve better ranking than single-elimination tournaments within same or less number of matches.
Appendix A Appendix
A.1 Machine Requirements for Experiments
Except the part we inferenced open-weight models such as Llama, Qwen and Gemma, our experiments are mostly do not require GPU usage. Inference are done on one A100 GPU, but T4 would be enough for reproducing our experiments. Otherwise, our experiments require querying API and post-processing those with CPU. Experiments could be run on personal desktops. The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
A.2 Assuring Statistical Significance of the Results within Budget for proprietary models
To ensure a statistically significant number of trials for each experiment while staying within budget, we utilize OpenAI’s Batch API to prepare full-grid match outcomes (i.e., all-play-all matches for every prompt) in a cache file, allowing us to reuse these outcomes. Each empirical experiment consists of 500 trials per setting, with results represented using whisker plots or summary statistics such as median values. When experimenting with a subset of the Arena-Hard-Auto benchmark ($|X|<500$), we sample a stratified subset of the benchmark dataset for each new trial.
Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al. (2024) prompts. VARCO Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category). Error bars are 95% confidence intervals.
A.3 Elo ratings from VARCO Arena compared to Human Annotations
Figure 4 shows the Elo ratings computed out of VARCO Arena. For judge, we used gpt-4o. As mentioned in the caption, the Elo ratings are bootstrapped median value from 500 trials. 95% confidence intervals also plotted as an error bar, which look negligible in scale compared to observed values. Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
A.4 Binary search vs. Win rate over reference
Algorithm 2 Binary Search for Enlisting new LLM to a leaderboard
1:Leaderboard $L$, new model $m_{\text{new}}$, test prompts $X$, outputs $O_{ij}$, assumes $|X|>|L|>n_{\text{comparisons}}$
2:Updated leaderboard $L^{\prime}$ with $m_{\text{new}}$ placed
3:$n_{\text{comparisons}}\leftarrow\lfloor\log_{2}(|L|)\rfloor$
4:$n_{\text{matches}}\leftarrow\lfloor|X|/n_{\text{comparisons}}\rfloor$
5:function BinarySearchPlacement($L,m_{\text{new}}$)
6:     X $\leftarrow$ Shuffle(X)
7:     X $\leftarrow$ concat(X;X)
8:     low $\leftarrow 0$
9:     high $\leftarrow|L|-1$
10:     while low $\leq$ high do
11:         mid $\leftarrow\lfloor(\text{low}+\text{high})/2\rfloor$
12:         wins $\leftarrow 0$
13:         for $i\leftarrow 1$ to $n_{\text{matches}}$ do
14:              $x\leftarrow$ $X$.pop()
[…]
16:                  wins $\leftarrow$ wins $+1$
17:              end if
18:         end for
19:         if wins $>n_{\text{matches}}/2$ then
20:              high $\leftarrow$ mid $-1$
21:         else if wins $<n_{\text{matches}}/2$ then
22:              low $\leftarrow$ mid $+1$
23:         else if $|X|>$then
24:              continue $\triangleright$ Ensure tie
25:         else
26:              return mid, tie $\triangleright$ Tie
27:         end if
28:     end while
29:     return low, non-tie $\triangleright$ Position found
30:end function
31:function UpdateLeaderboard($L,m_{\text{new}}$)
32:     position, istie $\leftarrow$ BinarySearchPlacement($L,m_{\text{new}}$)
33:     $L^{\prime}\leftarrow L$.insert(position, $m_{\text{new}}$, istie)
34:     return $L^{\prime}$
35:end function
A.4.1 Binary Search
We tried binary search placement of a newly added LLM to the leaderboard without reference text in Table 5. Details of how we implemented binary search are attached in Appendix 2. It turns out that binary search based on leaderboard ranks is not as reliable as the current approach of scoring the newcomer to the reference outputs. The number of judge operations performed is equivalent to the matches allocated to the least-performant model in a tournament, which is $|X|$ (i.e. maximum possible matches that an LLM could have is $|X|*\text{log}_{2}n_{\text{model}}$). Within the size of the benchmark prompts ($|X|$), binary search is incompatible with the current approach of using reference instead.
A.4.2 Comparing to the most performant Model so far: Converting Elo Table back to Win Rate
Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion. One could pick an appropriate anchor LLM as a reference in a leaderboard to estimate the skill of a newcomer. Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the reference. This is the same as converting the Elo table into a win rate leaderboard. Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the reference regardless we have picked. There are two ways to estimate the win rate over the reference model. We could just count the matches given are enough in amount, or we could also convert Elo ratings back to $P(i>a)$ to use it directly for scoring for the model ranks in the leaderboard. Reminding that Elo rating is purposed for expecting a likely outcome of the match, this should work. After this win rate of the newcomer model $P^{*}(n>a)=\frac{\text{count(n wins)}}{|X|}$ could be directly compared for enlisting.
$|\Delta_{\text{rank}}|$ ($\downarrow$) gt=1 2 3 4 5 6 avg.
binary search 0.09 1.24 1.75 1.55 1.26 1.10 0.92
(4o) (.04/-.03) (.14/-.14) (.09/-.09) (.07/-.06) (.08/-.08) (.10/-.09)
anchored 0.00 1.01 1.95 2.00 0.96 0.30 1.98
(4o) (0.00/0.00) (0.01/-0.01) (0.02/-0.02) (0.00/0.00) (0.02/-0.02) (0.04/-0.04)
binary search 0.52 0.85 0.59 2.03 1.20 2.45 1.27
(4o-mini) (.09/-.07) (.12/-.11) (.10/-.09) (.02/-.02) (.05/-.05) (.07/-.06)
anchored 0.00 0.00 1.00 2.00 2.00 1.00 1.00
(4o-mini) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00)
7 8 9 10 11 12 13 avg.
1.31 1.27 2.22 1.74 2.27 2.23 1.86 1.84
(.10/-.10) (.11/-.11) (.14/-.12) (.09/-.09) (.12/-.11) (.12/-.12) (.07/-.07)
0.30 3.68 1.09 1.03 2.97 0.78 1.00 1.55
(0.04/-0.04) (0.04/-0.04) (0.03/-0.03) (0.02/-0.01) (0.02/-0.02) (0.05/-0.05) (0.00/0.00)
0.69 0.85 3.89 1.95 2.10 2.37 0.88 1.82
(.07/-.06) (.09/-.09) (.12/-.11) (.06/-.05) (.03/-.03) (.10/-.11) (.12/-.11)
0.51 0.52 3.50 1.00 1.00 3.00 0.50 1.43
(0.49/-0.51) (0.48/-0.52) (0.49/-0.51) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.50/-0.50)
14 15 16 17 18 19 20 avg.
1.40 3.07 0.80 1.47 5.00 0.96 - 2.13
(.04/-.05) (.11/-.11) (.08/-.09) (.05/-.04) (.11/-.11) (.08/-.09)
2.00 2.00 1.00 1.21 3.00 0.21 - 1.57
(0.00/0.00) (0.00/0.00) (0.00/0.00) (0.03/-0.04) (0.00/0.00) (0.04/-0.03)
1.45 4.20 0.19 0.08 1.09 1.08 0.40 1.21
(.07/-.08) (.17/-.17) (.07/-.06) (.03/-.02) (.05/-.05) (.05/-.05) (.07/-.07)
1.00 2.00 2.00 1.00 1.00 3.00 0.00 1.43
(0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00)
Table 5: Binary search vs. Anchored comparison: Mean rank deviation ($|\Delta_{\text{rank}}|$) from ground-truth leaderboard. Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto Li et al. (2024)).
A.5 Separability In terms of Confidence Interval
To see how well the two benchmarking approach (anchored comparison and tournament approach) separates LLMs in adjacent ranks, we provide scatter plot of Elo rating and win rate paired with error bar (95% confidence interval). We present the both results of using gpt-4o (Figure 5) and gpt-4o-mini (Figure 5) as a judge. Inside the each plot, inseparables indicates the cases where any pair of datapoint co-includes each other within their range of error bars, and overlap means a certain datapoint is within some other’s range of error, when it is one-sided.
Figure 5: gpt-4o result of anchored comparison and tournament approach. 1000 bootstrapped median from 500 observations used for confidence interval estimation.
Figure 6: gpt-4o result of anchored comparison and tournament approach. 1000 bootstrapped median from 500 observations used for confidence interval estimation.
A.6 Judge configuration
A.6.1 Evaluation Prompt
We use the prompt from LLMBar. The prompt depicted in Figure A.6.2. We added 4 questions for criteria of our own to Metrics.txt prompt of Zeng et al. (2024). You can refer to the original prompt in LLMBar github.
A.6.2 Decoding Parameters
We did not configure decoding parameters of judge LLMs (gpt-4o[-mini]), which its temperature defaults to 1. The only parameter we have adjusted is maximum number of tokens to be generated, which for our prompt is less than 6 (i.e. The output of our prompt is (a) or (b)). To avoid position bias, we alternated the position of the responses from a certain model across the benchmark prompt.
PROMPTS = [ # metrics.txt from LLMBar
"role": "system", "content": "You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.", ,
"role": "user", "content": """Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
Here are some rules of the evaluation:
(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
(3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better.
Do NOT provide any explanation for your choice.
Do NOT say both / neither are good.
You should answer using ONLY "Output (a)" or "Output (b)". Do NOT output any other words.
# Instruction:
instruction
# Output (a):
response_a
# Output (b):
response_b
# Questions about Outputs:
Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
- Does the output well satisfy the intent of the user request?
- If applicable, is the output well-grounded in the given context information?
- Does the output itself satisfy the requirements of good writing in terms of:
1) Coherence
2) Logicality
3) Plausibility
4) Interestingness
# Which is better, Output (a) or Output (b)? Your response should be either "Output (a)" or "Output (b)":""",

### v4

@@
-VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models
+Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons
 Abstract
-Most existing benchmarking approaches for evaluating the output quality of large language models (LLMs) rely on comparing LLM responses to predefined references.
-Such methods, based on static datasets, quickly become outdated as LLM capabilities and use cases evolve.
-In this work, we introduce VARCO Arena—a novel, cost-effective, and robust benchmarking approach that leverages a single-elimination tournament structure to minimize the number of required comparisons while eliminating the need for static references or costly human annotations.
-We validate our approach through two experiments: (i) a simulation study that examines its robustness under various conditions, and (ii) an empirical evaluation using publicly available benchmark prompts.
-In both experiments, VARCO Arena consistently outperforms current LLM benchmarking practices, achieving stronger correlations with human-established Elo ratings.
-Our results demonstrate that VARCO Arena not only produces reliable LLM rankings but also provides a scalable, adaptable solution for qualitative evaluation across diverse, customized use cases.
-We release our demo and code at [https://huggingface.co/spaces/NCSOFT/VARCO_Arena].
-VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models
-Seonil Son1, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: deftson@ncsoft.com
+As Large Language Models (LLMs) expand across domains, LLM judges have become essential for systems evaluation.
+Current benchmarks typically compare system outputs against baselines.
+This baseline-mediated approach, though convenient, yields lower reliability than direct comparison between systems.
+We propose Arena-Lite which integrates tournament structure on top of head-to-head comparison.
+The application of a tournament structure and direct comparison eliminates the need for baseline outputs, reduces the number of required comparisons, and allows higher reliability in system rankings.
+We conducted two experiments: (1) controlled stochastic modeling and (2) empirical validation with a real LLM judge.
+Those experiments collectively demonstrate that Arena-Lite consistently achieves higher reliability with fewer comparisons, even with smaller datasets or weaker judges.
+We release an easy-to-use web demonstration and code to foster adoption of Arena-Lite, streamlining model selection across research and industry communities.
+Arena-Lite demo and code are available on https://huggingface.co/spaces/NCSOFT/ArenaLite
+Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons
+Seonil Son1thanks: Current institution: RLWRLD, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: seonil.calendar@gmail.com
 1 Introduction
-The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks.
-The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios Ouyang et al.
-(2022); Köpf et al.
+Figure 1: Arena-Lite directly compares LLM response pairs over multiple single-elimination tournaments rather than comparing responses to baseline outputs.
+In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
+Total no.
+matches ($\downarrow$)
+No. matches per
+LLM participant ($\uparrow$)
+[…]
+Arena-Lite (ours) $(n_{\text{model}}-1)\cdot|X|$ $[~|X|,~|X|*\lceil\log_{2}n_{\text{model}}\rceil~]$
+Table 1: Comparison between Current practice of benchmarking (comparing to baseline outputs) and Arena-Lite.
+$|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively.
+Arena-Lite, always save |X| number of comparisons for benchmarking, while allows more matches per LLM participant thanks to head-to-head comparison.
+LLMs excel in diverse tasks, from chatbots to code generation, due to their powerful generative capabilities Ouyang et al.
 (2022); Roziere et al.
 (2023).
-As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant.
-In response to this challenge, numerous benchmarks evaluating LLM capabilities have emerged Hendrycks et al.
+As their versatility grows, accurately evaluating their performance becomes critical.
+To address this, benchmarks like MMLU and BigBench have emerged to assess LLM capabilities across various domains Hendrycks et al.
 (2020); Srivastava et al.
-(2023); Zhong et al.
 (2023).
-Many LLM benchmarks employ formats amenable to automated scoring.
-Examples include benchmarks for arithmetic problems Gao et al.
-(2022); Patel et al.
-(2021), multiple-choice questions Lin et al.
-(2022), and code execution Austin et al.
+Many of these benchmarks, such as those for arithmetic or code execution (e.g., GSM-Hard, HumanEval Gao et al.
 (2022); Chen et al.
+(2021)), use automated scoring to evaluate problem-solving skills.
+However, their focus is not on quality of generated content or limited to the cases where the generated contents are automatically evaluated (e.g.
+programming), which are mostly not the case for variety of generation tasks.
+The Chatbot Arena, a leading platform for reliable human evaluation of LLMs, has set a standard by collecting extensive human annotations Chiang et al.
 (2024).
-While these benchmarks are valuable, they primarily assess problem-solving abilities.
-The need for LLM benchmarks extends beyond this, as the primary value of LLMs lies in the versatility of their generative behavior.
-Researchers have primarily relied on pairwise comparisons to evaluate and rank the LLMs based on human preference annotations.
-A representative example of this approach is Chatbot Arena Chiang et al.
-(2024), which computes Elo ratings based on a massive number of human votes.
-While Chatbot Arena benefited from reliable, open-ended prompts generated by a large user base, obtaining such a high volume of annotations from dedicated annotators remains challenging.
-Figure 1: VARCO Arena directly compares LLM response pairs in single-elimination tournament rather than comparing references.
-In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
-No. Comp. ($\downarrow$) Judge Eval. Type
-Chatbot Arena unknown human head-to-head
-Current Practice $n_{\text{model}}\cdot|X|$ LLM reference -based
-VARCO Arena $(n_{\text{model}}-1)\cdot|X|$ LLM head-to-head
-Table 1: Comparison between Current Practice and Varco Arena.
-$|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively.
-Human annotators are considered much more costly than LLM judge counterpart.
-To address these challenges, several benchmark datasets for generation tasks have been developed and applied with an LLM judge Zheng et al.
-(2024) to quantify LLM capabilities.
-Using these benchmarks and their reference responses, LLMs are ranked via pairwise comparisons between their outputs and the reference responses provided by LLM judges.
-Representative examples include AlpacaEval Li et al.
-(2023), Arena-Hard-Auto Li et al.
-(2024), and MTBench Zheng et al.
+Yet, its resource-intensive approach has prompted efforts to replicate its rankings using LLM judges as a cost-effective alternative Li et al.
+(2024, 2023).
+These methods, however, rely on baseline-mediated comparisons—comparing LLM outputs to a baseline encore model’s outputs—which sacrifice reliability.
+Current benchmarks relying on encore models often rank LLMs by their win rate against baseline responses from an encore model.
+This approach has two advantages: it scales linearly with the number of LLMs and provides a consistent quality standard.
+However, we argue that comparing LLMs directly against each other is inherently more reliable than using baseline outputs, which can introduce noise coming from weak transitivity Xu et al.
+(2025) of human preferences on LLM responses.
+To address this, we propose Arena-Lite, a novel evaluation framework that uses direct, head-to-head comparisons organized in a tournament structure.
+By eliminating the need for baseline outputs, Arena-Lite reduces the number of comparisons required while achieving stronger alignment with human-established rankings, such as those from Chatbot Arena.
+Arena-Lite conducts single-elimination tournaments among participating LLMs for each prompt.
+From the match results, we can compute Bradley-Terry preference ratings for the final ranking Bradley and Terry (1952).
+This results in a single scalar per model that captures relative performance between any model counterpart, enabling accurate and efficient ranking.
+We validate Arena-Lite through two experiments.
+The first experiment, stochastic modeling of LLM competition (Section 4.2) demonstrates that tournament-based direct comparison method outperforms baseline-mediated method under various conditions, including different numbers of LLM participants, dataset rows used, and judge accuracies.
+Second, our empirical experiment (Section 4.3) shows that Arena-Lite achieves higher correlation with Chatbot Arena’s rankings than standard approaches using baseline outputs (Table 1) attested toward number of LLM as judges.
+These results collectively highlight Arena-Lite’s ability to deliver reliable rankings with fewer comparisons, even with smaller datasets or weaker judges over various generation tasks.
+Our contributions are threefold:
+  1. 1.
+    We introduce Arena-Lite, a tournament-based framework for direct LLM comparisons, offering greater reliability than baseline-mediated approaches.
+  2. 2.
+    We rigorously demonstrate, through both comprehensive modeling and empirical experiments, that Arena-Lite achieves more accurate LLM rankings while requiring fewer comparisons than prevalent practices, particularly those relying on common baseline model outputs.
+  3. 3.
+    We open-source a functional demo and the complete code for Arena-Lite (https://huggingface.co/spaces/NCSOFT/ArenaLite), enabling researchers and industry practitioners to easily host and utilize our framework for streamlined LLM evaluation.
+2 Preliminaries: Quantifying Generative Performance
+Quantifying the generative capabilities of LLMs is an inherently challenging task.
+The evaluation is complicated by the stochasticity of model outputs and the subjectivity of human judgments.
+To approximate real-world performance, a common methodology involves assessing model outputs across a diverse range of prompts.
+In this context, two metrics are widely utilized: the win rate, which measures the frequency of preference for a model’s output over a baseline, and the Bradley-Terry model, which is employed to infer a latent skill rating for each model based on pairwise comparisons.
+2.1 Measuring Win rate over baseline outputs
+Benchmarks like AlpacaEval and Arena-Hard-Auto assess LLM response quality by comparing it to baseline responses from a encore model Li et al.
+(2023, 2024).
+An LLM judge evaluates whether the candidate LLM’s response outperforms the baseline for a given prompt.
+The win rate—the proportion of prompts where the LLM’s response is preferred—serves as a measure of its generative ability.
+While this approach is straightforward and scalable, it introduces noise coming from mediated comparisons.
+2.2 Bradley-Terry Model Preference for LLM Rating
+The Bradley-Terry (BT) model Bradley and Terry (1952) is widely used to infer baseline-mediated rankings of LLMs from pairwise comparisons.
+Chatbot Arena adopts the BT model rather than the classical Elo system Elo and Sloan (1978), but both Elo and BT models are useful for expecting probability of match outcome based on a score difference, though they differ in update rules and statistical assumptions.
+In the BT model, each LLM is assigned a latent score representing its procificency.
+Given LLMs $i$ and $j$ with scores $R_{i}$ and $R_{j}$, respectively, the probability that LLM $i$ is preferred over LLM $j$ is modeled as:
+$\text{P}(i>j)=\frac{1}{1+10^{(R_{j}-R_{i})/400}}.$ (1)
+This formulation closely resembles the Elo win-probability function, reinforcing the intuitive connection between the two.
+Chatbot Arena uses this BT-based formulation to rank LLMs by aggregating human preferences collected through pairwise matchups Chiang et al.
 (2024).
-There are two major advantages to relying on reference responses: (1) The number of comparisons required for ranking scales linearly with the number of LLMs.
-(2) The reference response establishes a quality standard for evaluating LLM outputs.
-However, we argue that using reference responses to mediate comparisons between LLM outputs is suboptimal and that head-to-head comparisons yield more accurate assessments.
-In this context, we propose VARCO Arena as a novel evaluation framework that directly compares LLM responses while requiring fewer comparisons.
-Even with fewer comparisons and without relying on reference responses, VARCO Arena achieves a stronger correlation with the human-established Elo rankings from Chatbot Arena.
-VARCO Arena is designed to conduct a single-elimination tournament for each prompt across all participating LLM responses and then compute Elo ratings Elo and Sloan (1978) based on the outcomes of these matches.
-By organizing tournament-style matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs with fewer comparisons than those required by reference-based methods.
-These win rates are subsequently interpreted as Elo ratings, enabling us to generate a comprehensive ranking of candidate LLMs.
-We validate the effectiveness of VARCO Arena against current LLM benchmarking practices through two experiments.
-First, we conduct a simulation study (Section 4.2) to evaluate the reliability and robustness of our tournament approach under various controlled factors, including the number of participating LLMs, the number of test prompts, and the accuracy of the judge model.
-Second, we empirically demonstrate that running VARCO Arena with several judge LLMs on public benchmark (Section 4.3) prompts yields a stronger correlation with human-established Elo rankings from Chatbot Arena than does the reference-based approach.
-These experiments confirm that our benchmarking approach outperforms current practices in terms of ranking reliability while requiring the lesser number of comparisons (Table 1).
-Notably, the simulation results—which emulate outcomes under various conditions—underscore that the superiority of our tournament approach is not coincidental but rather reflects the inherent advantages of direct pairwise comparisons over mediated comparisons using reference responses.
-In this work, we make the following contributions:
-  1. 1.
-    We introduce VARCO Arena, a novel reference-free benchmarking method for LLMs that leverages a single-elimination tournament structure to perform head-to-head comparisons of LLM responses across test prompts.
-This approach eliminates the dependency on static references and enhances benchmarking flexibility.
-  2. 2.
-    Through extensive simulation studies and empirical evaluations, we demonstrate that VARCO Arena not only yields more reliable LLM rankings compared to traditional reference-based methods but also achieves this with significantly fewer comparisons, thereby reducing evaluation costs.
-  3. 3.
-    We release our demo and code to facilitate the further research and reproducibility; our resources are available at [https://huggingface.co/spaces/NCSOFT/VARCO_Arena]
-2 Preliminaries: Quantifying Generation Ability
-Quantifying an LLM’s generation ability is crucial, but it presents several challenges.
-The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as the provided test prompts and the inherently subjective nature of human preferences.
-A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its real-world performance.
-Two popular measures for LLM generation ability are the win rate against reference responses and the Elo rating.
-2.1 Win Rate in Comparison to the Reference
-AlpacaEval Li et al.
-(2023) and Arena-Hard-Auto Li et al.
-(2024) are representative benchmark datasets that aim to quantify LLM response quality using reference responses.
-These benchmarks employ an LLM judge with a standardized prompt for automated evaluation of generation capability.
-Given the prompts, LLM responses, and reference responses, the judge LLM is tasked with determining whether the reference or the response is preferred.
-The LLM’s win rate across the test prompts in the benchmark is used as a measure of its generation proficiency.
-2.2 Elo Rating
-The Elo rating system, introduced by Elo and Sloan, has since become a popular method for quantifying performance levels in competitive sports and, more recently, for evaluating the generative abilities of LLMs.
-The primary purpose of the Elo rating system is to represent a participant’s skill level as a single scalar value, enabling the prediction of relative win rates between participants who have never directly competed.
-Assessing the superiority of one LLM over another in terms of generative capability shares several similarities with determining winners in competitive sports.
-Varying test prompts may yield different results, similar to how weather or other factors can influence the outcome of a sports match.
-While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently.
-With Elo ratings, one can expect the relative chance of winning between a pair of players.
-The expected win rate of the player $i$ against $j$ is computed as follows:
+Users are shown responses from two anonymized models to the same prompt and asked to select which response they prefer.
+The accumulated judgments are then used to fit BT scores, producing a leaderboard that reflects relative model performance.
+While this approach requires a substantial number of human evaluations to ensure reliability, it captures nuanced quality differences between models more effectively than purely automatic benchmarks.
+Arena-Lite, introduced in the next section, builds on the same BT modeling framework but seeks to reduce the number of required comparisons by using tournament-structured match-making.
+3 Arena-Lite
+To address the high annotation cost of Chatbot Arena while preserving evaluation reliability, we propose Arena-Lite.
+Arena-Lite introduces a tournament-based approach for efficient and reliable LLM evaluation using a single-elimination structure.
+Unlike baseline-mediated evaluations that compare model outputs to a baseline, Arena-Lite directly compares outputs from different models through head-to-head matchups for each prompt in benchmark datasets.
+Repeated tournaments across the dataset produce consistent leaderboards reflecting models’ fundamental performance.
+We first discuss limitations of baseline-mediated evaluations (Section 3.1).
+Next, we describe how Arena-Lite conducts tournaments to generate ratings (Section 3.2, Algorithm 1).
+Finally, we highlight similarities between the single-elimination structure and merge sort, explaining why aggregated tournaments yield reliable LLM rankings (Section 3.3).
+3.1 Comparing to Baseline outputs is not Always Helpful
+Although baseline outputs are a standard way to evaluate and rank LLMs, they introduce potential failure modes.
+Beyond the fact that a single baseline output might not capture every dimension of appropriate answers, relying solely on a baseline output can lead to unreliable rankings of LLMs.
+Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs.
+If we choose to compare LLM responses directly to rank them using BT preference (Equation 1), all head-to-head comparisons are utilized.
+In contrast, baseline-mediated evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
+$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ \quad\text{vs.}\,\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\hfil\hskip 20.00003pt&\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})>Y_{i}&\text{(unhelpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})<Y_{i}&\text{(unhelpful)}\end{cases}\end{array}$
+When the baseline output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the baseline is effective for benchmarking.
+Otherwise, these comparisons do not help differentiate LLM performance.
+Consequently, the baseline-mediated approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the baseline.
+3.2 Tournaments of LLMs over multiple prompts to preference ratings
+Algorithm 1 Tournament-Based Model Evaluation
+1:Models $M=\{m_{1},m_{2},\ldots,m_{n}\}$, Prompts $X=\{x_{1},x_{2},\ldots,x_{k}\}$
+2:Bradley-Terry preference ratings
+3:Initialize $R\leftarrow\emptyset$
+4:for each $x_{j}\in X$ do
 […]
-Computing accurate Elo ratings requires a sufficient number of matches to estimate relative win rates between pairs of participants.
-The resulting matrix of relative win rates is then used to estimate Elo ratings through logistic regression, similar to multivariate logistic regression with a sigmoid function.
-In Equation 1, $P(i>j)$ represents the expected win rate of participant $i$ over $j$, $R_{i}$ is the Elo rating of participant $i$, and $\Delta_{\text{ij}}$ is the Elo rating difference between participants $j$ and $i$ ($R_{j}-R_{i}$).
-Chatbot Arena presents a leaderboard of LLMs, with Elo ratings computed from matches evaluated by a large user base.
-Users prompt a pair of LLMs and submit their judgments of which one responded better.
-Although Chatbot Arena relies on manual evaluation of matches, it offers an intuitive method for comparing LLMs using Elo ratings.
-3 VARCO Arena
-We propose VARCO Arena, a reference-free approach for benchmarking large language models (LLMs) via single-elimination tournaments.
-Instead of comparing model responses against fixed references, VARCO Arena directly compares outputs from different models, determining superiority through head-to-head matchups for each prompt in our benchmark dataset.
-Repeated tournaments across prompts yield reliable leaderboards that reflect the relative performance of each model.
-We begin by motivating our approach over current reference-based evaluations (Section 3.1).
-Next, we detail how VARCO Arena performs tournaments and results in Elo ratings (Section 3.2 and Algorithm 1).
-Finally, we discuss the reason why aggregating the results from tournaments promises reliable ranking (Section 3.3).
-3.1 Comparing to a Reference is not Always Helpful
-Although reference texts are a standard way to evaluate and rank large language models (LLMs), they introduce potential failure modes.
-Beyond the fact that a single reference might not capture every dimension of correctness, relying solely on a reference can lead to unreliable rankings of LLMs.
-Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs.
-If we choose to compare LLM responses directly to rank them using Elo ratings (Equation 1), all head-to-head comparisons are utilized.
-In contrast, reference-based evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
-$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ \quad\text{vs.}\,\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\hfil\hskip 20.00003pt&\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})>Y_{i}&\text{(unhelpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})<Y_{i}&\text{(unhelpful)}\end{cases}\end{array}$
-When the reference output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the reference is effective for benchmarking.
-Otherwise, these comparisons do not help differentiate LLM performance.
-Consequently, the reference-based approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the reference.
-3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
-Algorithm 1 Tournaments of LLMs over prompts
-1:prompts $X=\{x_{1},~{}x_{2},~{}...,~{}x_{i}\}$, LLMs $M=\{m_{1},~{}m_{2},...,~{}m_{j}\}$, outputs $O_{i,j}$ = $m_{j}(x_{i})$
-2:Ranked LLMs with Elo ratings
-3:function Match($m_{1},m_{2},x$)
-4:     return $m_{1}$ if $\texttt{IsBetter}(O_{x,1},O_{x,2})$
-5:     else $m_{2}$
-6:end function
-7:function SingleElim($M$, $x$, res)
-8:     if $|M|=2$ then
-9:         res.append(Match($M$[0], $M$[1], $x$))
-10:         return res[-1]
-11:     end if
-12:     mid $\leftarrow\lfloor|M|/2\rfloor$
-13:     left $\leftarrow$ SingleElim($M$[:mid], $x$, res)
-14:     right $\leftarrow$ SingleElim($M$[mid:], $x$, res)
-15:     return SingleElim(left + right, $x$, res)
-16:end function
-17:function Tournaments2Ranks($X,M$)
 […]
-19:     for $x_{i}\in X$ do
-20:         SingleElim(Shuffled($M$), $x_{i}$, res)
-21:     end for
-22:     return ComputeElo(res)
-23:end function
-Figure 1 and Algorithm 1 illustrate how VARCO Arena benchmarks LLMs via a tournament approach.
+7:  $M^{\prime}\leftarrow M\cup\{\texttt{None}\}^{\texttt{n\_byes}}$
+8:  Randomly shuffle $M^{\prime}$
+[…]
+10:end for
+[…]
+[…]
+13:  if $|M^{\prime}|=2$ then
+[…]
+15:   if $\texttt{None}\notin M^{\prime}$ then
+16:     Add result to $R$
+17:   end if
+[…]
+[…]
+20:   end if
+21:  else
+22:   $\texttt{mid}\leftarrow\lfloor|M^{\prime}|/2\rfloor$
+[…]
+[…]
+25:  end if
+26:end function
+[…]
+28:  if $m_{i}=\texttt{None}$ and $m_{j}=\texttt{None}$ then
+29:   return None
+30:  else if $m_{i}=\texttt{None}$ or $m_{j}=\texttt{None}$ then
+31:   return $m_{i}$ if $m_{j}=\texttt{None}$ else $m_{j}$
+32:  else
+33:   $O_{i}\leftarrow m_{i}(x)$, $O_{j}\leftarrow m_{j}(x)$
+34:   if $O_{j}>O_{i}$ then return $(m_{j},m_{i})$
+35:   elsereturn $(m_{i},m_{j})$
+36$\triangleright$ returns (winner, loser) tuple
+37:   end if
+38:  end if
+39:end function
+Figure 1 and Algorithm 1 illustrate how Arena-Lite benchmarks LLMs via a tournament approach.
 Here, $|X|$ denotes the number of prompts in the benchmark dataset.
-Each execution of VARCO Arena runs a tournament among participant LLMs for every prompt in the dataset.
+Running Arena-Lite hosts tournaments among participant LLMs for every prompt in the dataset.
 The use of tournament structures for LLM benchmarking offers both benefits and challenges.
 A major advantage of a single-elimination tournament is efficiency.
-As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using references.
+As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using baseline outputs.
 However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
@@
   3. 3.
-    Aggregating match results allows for precise win rate estimation via Elo ratings, resulting in a well-aligned overall ranking.
+    Aggregating match results allows for precise win rate estimation via BT preference, resulting in a well-aligned overall ranking.
   4. 4.
-    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt.
-In Section 3.3, we further explain how aggregating multiple tournaments could yield an reliable ranking of LLMs.
-We also provide an analysis of the number of matches each LLM faces, offering a comprehensive view of the method’s efficiency and effectiveness.
-3.3 Why Aggregating Multiple Tournaments Yields Reliable Ranks?
-Our approach aggregates match outcomes from multiple tournaments to approximate the complete set of pairwise comparisons—akin to the comparisons made in merge sort.
-In a single-elimination tournament, every participant advances based solely on match outcomes, a process that mirrors the merging steps in merge sort.
-Notably, a single-elimination tournament executes only the comparisons strictly necessary for determining the winner, omitting many comparisons that merge sort would perform further.
-We posit that the missing pairwise match-ups in any one tournament can be recovered by aggregating tournaments conducted over different prompts.
-This hypothesis relies on the assumption—central to our use of the Elo model—that match outcomes are independent of the prompt.
-Consequently, matches across different prompts are considered equivalent.
-Considering only the initial comparisons, which are randomly sampled, the aggregate number of comparisons is $|X|\cdot n_{\text{model}}/2$.
-Since $n_{\text{model}}$ is typically on the order of tens and $|X|$ comprises at least hundreds, this number exceeds the total possible match-ups, $\binom{n_{\text{model}}}{2}$ [^1] Note that merge sort produces a set of match-ups with no duplicates..
-Furthermore, as shown in Table 1, the remaining matches—totaling $|X|\cdot(n_{\text{model}}-1)$—either contribute additional merge sort structures or enhance the accuracy of estimating relative win rates among LLM participants.
-Moreover, considering that each unique pair meets in at least $|X|/(2(n_{\text{model}}-1))$ matches [^2] A model participates in between $|X|$ and $\lceil\log_{2}n_{\text{model}}\rceil$ matches per tournament.
-Dividing this by the number of possible match-ups per model, $2(n_{\text{model}}-1)$, yields the expression above.
-across the benchmark, this frequency is sufficient for a fair estimation of their relative win rates.
-In summary, aggregating tournaments not only reconstructs the full set of pairwise comparisons for a merge sort but also ensures that each pair of models faces one another often enough to yield accurate win rate estimations, leading to reliable Elo ratings.
-Based on these considerations, we propose that conducting tournaments over a benchmark prompts will yield a reliable ranking of LLMs.
+    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt (Table 1).
+In Section 3.3, we further explain how aggregating multiple tournaments could yield reliable ranking of LLMs.
+We also provide further analysis on number of comparisons performed over tournaments of Arena-Lite comparing to merge sort, offering a comprehensive view of the method’s efficiency and effectiveness.
+3.3 Why Aggregating Multiple Tournaments Yields Reliable Rankings
+[…]
+To achieve reliable rankings of LLMs, our approach aggregates match outcomes from multiple tournaments, lesser than a full grid comparisons, but effectively approximating the complete set of pairwise comparisons required for merge sort.
+We outline the rationale in four key points:
+Merge Sort Baseline A single-elimination tournament mirrors the merging steps of merge sort, which requires $\mathcal{O}(n\log n)$ comparisons with no duplicate match-ups to rank $n$ models.
+However, a single tournament omits many comparisons, covering only the minimal match-ups needed to determine a winner.
+Recovering Comparisons via Aggregation By aggregating tournaments over diverse prompts, we can recover missed pairwise match-ups that should have occurred.
+Assuming match outcomes are prompt-independent (as BT model assumes), matches across prompts are considered equivalent.
+With $|X|$ prompts (typically hundreds to thousands) and $n_{\text{model}}$ models (tens), only random initial match-ups totaling $|X|\cdot\lfloor\frac{n_{\text{model}}}{2}\rfloor$.
+This exceeds the $\binom{n_{\text{model}}}{2}$ possible combination, ensuring broad coverage.
+Sufficiency of Comparisons The aggregated match-ups not only cover the necessary comparisons but also surpass the $\mathcal{O}(n\log n)$ requirement of merge sort.
+Moreover, each unique model pair competes in average $|X|/(n_{\text{model}}-1)$ to $\frac{|X|\log_{2}n_{\text{model}}}{n_{\text{model}}-1}$ matches [^1] These estimates are computed by number of matches a model undergoes which is $[~|X|,~|X|*\lceil\log_{2}n_{\text{model}}\rceil~]$, devided by number of possible opponents for a model, $n_{\text{model}}-1$ across the benchmark, a frequency mostly sufficient for accurate win rate estimation.
+Refinement for Reliability The remaining matches, totaling $|X|\cdot(n_{\text{model}}-1)$, further refine the ranking by enhancing win rate estimates, especially among top-performing models, reducing noise and ensuring robustness akin to Arena-Lite’s sampling strategy.
+In summary, aggregating multiple tournaments reconstructs the full set of comparisons needed for a merge sort-like ranking while providing enough repeated match-ups to ensure accurate win rate estimations.
+This dual mechanism yields reliable and robust LLM rankings across the benchmark.
 4 Experiments
-We run two experiments with different settings for comparing VARCO Arena and current practice of reference-based benchmarking.
-In the first experiment (Section 4.2) we conduct a simulation study to control various factors that affect LLM benchmarking.
-This simulation tests our foundational propositions for VARCO Arena design (mentioned in Section 3.1 and 3.3) under a more controlled, simplified environment immune to noisy factors such as potential biases of LLM judges Park et al.
+We conducted two experiments to evaluate Arena-Lite against baseline-mediated benchmarking.
+The first experiment (Section 4.2) utilized a stochastic model to simulate LLM competitions, comparing Arena-Lite’s tournament-based direct comparison with baseline-mediated evaluation.
+This controlled setup allowed us to test Arena-Lite’s design principles, such as the effectiveness of direct versus mediated comparison (Section 3.1) and tournament-based sampling (3.3), while isolating variables and minimizing noise, such as LLM judge biases Park et al.
 (2024).
-The other is empirical experiments (Section 4.3).
-We use gpt-4o and -mini as well as other several popular models such as Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to validate the effectiveness of our tournament approach against current LLM benchmarking practices.
-By presenting both simulation and empirical results, we aim to demonstrate the effectiveness of our tournament approach.
-In Section 4.1, we first describe the common experimental settings before getting into specific details of each experiments in the following subsections.
-4.1 Chatbot Arena Leaderboard Ratings as Ground-Truth LLM Rankings
-We compare the results of each benchmarking approach against the rankings from the Chatbot Arena leaderboard.
-Chatbot Arena is widely regarded as one of the most reliable leaderboards due to its extensive collection of human preference annotations.
-Given the large number of votes and the diverse set of prompts used for model comparisons, the resulting rankings are considered sufficiently accurate to serve as ground truth.
-4.2 Experiment 1: Simulation Study
-We designed a simple simulation to emulate a probabilistic model of LLM matches, adhering to the Elo rating system in a controlled environment.
-In line with the Elo model’s assumptions, our judge is configured to follow Equation 3 exactly.
-The judge stochastically determines the winner of each LLM match solely based on the Elo rating difference ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$).
-As described in Equation 3, the outcome of a single match is sampled according to $P_{\text{predict}}(i>j)$, which is computed as the product of the judge’s accuracy and the likelihood of model $i$ beating model $j$ based on the Elo gap ($P_{gt}$).
+The second experiment (Section 4.3) validates Arena-Lite empirically using various LLMs as judges and public benchmark data.
+We tested models including gpt-4o, gpt-4o-mini, Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to assess Arena-Lite’s effectiveness against standard benchmarking practices.
+Together, these experiments demonstrate the superior reliability and efficiency of Arena-Lite’s tournament approach.
+Section 4.1 outlines shared experimental settings, followed by detailed descriptions of each experiment in subsequent subsections.
+4.1 Chatbot Arena Leaderboard as Ground-Truth Rankings
+We benchmark Arena-Lite and baseline-mediated evaluation against rankings from the Chatbot Arena leaderboard, widely recognized for its reliability due to extensive human preference annotations.
+With a large volume of votes across diverse prompts, these rankings provide a robust ground truth for model comparisons.
+4.2 Experiment 1: Controlled Stochastic Modeling of LLM Competitions
+We suggest a simple stochastic model based on the Bradley-Terry (BT) framework to compare Arena-Lite’s approach with baseline-mediated evaluation.
+This experiments simulates prompt-agnostic LLM competitions which follows the Bradley-Terry model’s presumption, with outcomes determined by a judge following Equation 3.
+The judge’s decision is based on the BT preference difference ($\Delta_{\text{ij}}$) between models $i$ and $j$, and the judge’s accuracy ($P_{\text{judge}}$):
 $\begin{split}P_{\text{predict}}(i>j)&=P_{\text{judge}}\times P_{\text{gt}}(i>j)\\ &=P_{\text{judge}}\times\frac{1}{1+10^{\Delta_{\text{ij}}/400}}\end{split}$ (3)
-The details of our simulation settings are as follows:
-Ground-truth Elo ratings (initial parameter of the simulation, and at the same time, the gold ranking to reproduce): We extracted Elo ratings from the English category of Chatbot Arena as of June 23.
-This Elo ratings are the estimates from massive user-submitted judgments (approximately 60% of the total submissions to the platform).
-For the simulation, any set of Elo ratings could be used, but we opted for real values computed from human preferences.
-Judge Accuracy ($P_{\text{judge}}$): In practice, a judge’s accuracy is an adaptive value that depends on factors such as the specific prompt-response pair and the manner in which LLM judges are prompted.
-In this simulation study, we control this parameter, varying it from 0.6 to 0.9.
-Number of Participant LLMs ($n_{\text{model}}$) and Benchmark Dataset Size ($|X|$): We varied these factors to assess the robustness of both the tournament and reference-based approaches under diverse conditions.
-This allowed us to explore how reliability changes with the number of participants in both data-poor and data-rich environments.
+With the model of judge above (Equation 3), we simulate both Arena-Lite’s tournament-based approach and baseline-mediated approaches according to the following initial conditions and procedures.
+Initial conditions:
+  • Ground-Truth BT Preference: We extracted BT preferences from the English category of Chatbot Arena (as of June 23), derived from approximately 60% of user-submitted judgments.
+These preferences serve as both the initial model parameters and the ground-truth rankings for evaluation.
+  • Judge Accuracy ($P_{\text{judge}}$): We varied judge accuracy from 0.6 to 0.9.
+[…]
 Simulation Procedure:
   1. 1.
-    Select the participant LLMs and obtain their Elo ratings for the simulation.
+    Select participant LLMs and their BT preferences.
   2. 2.
-    Compute the expected relative win rates ($P_{\text{gt}}$, see Equation 3) from the participants’ Elo ratings.
+    Compute expected win rates ($P_{\text{gt}}$) using Equation 3.
   3. 3.
-    Sample match outcomes for each LLM pair according to the benchmarking approach.
-The winner of each match is sampled from $P_{\text{pred}}$ (Equation 3), which depends solely on the Elo gap ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$).
+    Sample match outcomes based on $P_{\text{predict}}$ (Equation 3), determined by the BT preference rating gap ($\Delta_{\text{ij}}$) and judge accuracy ($P_{\text{judge}}$).
   4. 4.
-    Repeat step 3 for the designated number of test prompts ($|X|$).
+    Repeat for the specified number of test prompts ($|X|$).
   5. 5.
-    Calculate scores for ranking:
-    1. (a)
-      For the reference-based approach (current practice), use the win rate against the reference model (gpt-4-1106-preview, which has an Elo rating of 1233).
-    2. (b)
-      For the tournament approach (VARCO Arena), compute Elo ratings from all match outcomes.
+    Compute scores:
+    • Baseline-mediated: Win rate against a reference model (gpt-4-1106-preview, rating 1233).
+    • Arena-Lite: BT preference from all tournament match outcomes.
   6. 6.
-    Rank the models based on these scores.
+    Rank models based on scores.
   7. 7.
-    Compute the Spearman correlation between the simulated rankings (from step 6) and the ground-truth rankings (from step 1).
-We perform 50 trials for each simulation configuration to mitigate randomness from tournament brackets and sampling.
-4.3 Experiment 2: Assessing VARCO Arena Using Various LLM Judges
-To empirically validate our proposal, we evaluated the reliability of both VARCO Arena and reference-based approach over the top 20 models from the Chatbot Arena leaderboards.
+    Calculate Spearman correlation between simulated and ground-truth rankings.
+We conducted 50 trials per configuration to account for stochasticity in initial tournament brackets and judging process.
+4.3 Experiment 2: Empirical Validation of Arena-Lite with real LLM Judge
+To empirically validate our proposal, we evaluated the reliability of both Arena-Lite and baseline-mediated approach over the top 19 models from the Chatbot Arena leaderboards.
 This experiment employs actual prompt inputs and LLM outputs, distinguishing it from the earlier simulation study.
@@
 (2023), which comprises 800 prompt-reference pairs, could serve as a viable testbed, we opted for Arena-Hard-Auto because its design aligns more closely with Chatbot Arena.
-Arena-Hard-Auto uses responses from gpt-4-0314 as the reference outputs.
-For ranking, we utilized the reserved outputs of the top 21 models from the Arena-Hard-Auto Browser.
+Arena-Hard-Auto uses responses from gpt-4-0314 as the baseline outputs.
+For ranking, we utilized the reserved outputs of the top 20 (=19 + baseline) models from the Arena-Hard-Auto Browser.
 [^2] Extracted from the 2024 Jul 6 commit (fd42026).
 4.3.2 Participant LLMs
-For ranking, we selected 20 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
+For ranking, we selected 19 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
 4.3.3 LLM Judges
@@
 (2024) (See Appendix A.8.2).
-The same judge prompt was applied consistently across both the tournament and reference-based approaches.
+The same judge prompt was applied consistently across both the tournament and baseline-mediated approaches.
 To mitigate position bias Wu and Aji (2023), the order of model responses was alternated during evaluation.
@@
 The two experimental settings are summarized as follows:
-Experiment 1 (Simulation Study): This experiment uses the ground truth Elo ratings of the models to initialize the simulation.
-We vary control parameters for the benchmarking approaches—including the judge’s accuracy ($P_{\text{judge}}$), the number of test prompts used ($|X|$), and the number of participant LLMs ($n_{\text{model}}$)—to determine which benchmarking approach more accurately reproduces the participants’ ranking.
-For each configuration, we conduct 50 trials of experiments.
-Experiment 2 (Empirical Runs): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges.
-We select the top 20 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts.
-For both the tournament and reference-based approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings.
+[…]
+Experiment 2 (Empirical Validation): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges.
+We select the top 19 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts.
+For both the tournament and baseline-mediated approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings.
 In our empirical study, we conduct 500 trials for each experimental setting.
-Figure 2: Simulation results comparing the tournament and reference-based approaches.
-The tournament method consistently outperforms the reference-based approach in Spearman correlation across various control variables: the number of participant LLMs ($n_{\text{models}}$), the number of benchmark prompts ($|X|$), and judge precision ($P_{\text{judge}}$).
+Figure 2: Comparison of LLM ranking reliability between Arena-Lite and a baseline method in a stochastic simulation (Experiment 1, Sec.
+4.2).
+Ranking reliability is measured by the Spearman correlation ($\uparrow$) between the competition-derived ranking and the ground-truth ranking.
+Each box plot summarizes the results from 50 trials.
+The subplots analyze the effect of varying (from left to right) the number of competing models ($n_{\text{models}}$), the number of prompts ($n_{\text{prompts}}$), and the accuracy of the judge ($P_{\text{judge}}$).
+The single-elimination structure of Arena-Lite results in consistently higher correlation scores.
 5 Results and Discussion
-We assess the reliability and robustness of VARCO Arena as a means for LLM benchmarking, comparing it against the current reference-based approach.
-Our results from both simulation study and empirical runs indicate that the tournament approach of VARCO Arena yields rankings that align more closely with the ground-truth Elo leaderboards.
+We assess the reliability and robustness of Arena-Lite as a means for LLM benchmarking, comparing it against the current baseline-mediated approach.
+Our results from both simulation study and empirical runs indicate that the tournament approach of Arena-Lite yields rankings that align more closely with the ground-truth Chatbot Arena leaderboards.
 We present our findings using whisker plots and tables in the following sections.
-5.1 Experiment 1: Simulation Study Results
-Figure 2 illustrates noticeable differences in Spearman correlation, indicating that the tournament approach is more reliable than the reference-based method.
-The consistent performance gap across various conditions—namely, the number of participants, the number of test prompts, and judge accuracy ($n_{\text{model}}$, $|X|$, and $P_{\text{judge}}$)—demonstrates the robustness of the tournament approach.
-Although the simulation simplifies real-world complexity, a similar performance gap was observed in the empirical findings (Experiment 2, Figure 3).
-This consistency suggests that the robust performance of VARCO Arena is not coincidental or limited to a specific empirical setting of ours.
-5.2 Experiment 2: Empirical Results
-Figure 3: Results of VARCO Arena (tournament) and reference-based approach with gpt-4o (left) and gpt-4o-mini (right) judge.
-VARCO Arena constantly records higher Spearman correlation coherent with the Experiment 1 result (Figure 2).
-Results summary is on Table 2.
-As hinted in the previous section, the empirical results in Figure 3 show that VARCO Arena consistently outperforms the reference-based approach.
+5.1 Experiment 1: Modeling Experiment Results
+[…]
+5.2 Experiment 2: Empirical Validation Results
+Figure 3: Ranking reliability of Arena-Lite vs.
+utilizing baseline outputs.
+Arena-Lite consistently demonstrates higher Spearman’s rank correlation across numbers of benchmark prompts ($|X|$), indicating more reliable ranking.
+The evaluation was performed using gpt-4o (left) and gpt-4o-mini (right) as judge models, with a fixed number of models ($n_{\text{models}}$=19).
+Each box plot summarizes the results of 50 runs.
+(Experiment 2, Sec.
+4.3).
+As hinted in the previous section, the empirical results in Figure 3 show that Arena-Lite consistently outperforms the baseline-mediated approach.
 Although the performance gaps are less pronounced than in the simulation, the same trend persists.
-In Table 2, we report the median values for VARCO Arena and the reference-based approach using the gpt-4o family of judges while varying the number of test prompts ($|X|$).
-These results consistently demonstrate that VARCO Arena outperforms the reference-based method.
-Note that VARCO Arena shows similar or superior reliability even in extreme data-poor benchmark condition ($|X|=50$).
+In Table 2, we report the median values for Arena-Lite and the baseline-mediated approach using the gpt-4o family of judges while varying the number of test prompts ($|X|$).
+These results consistently demonstrate that Arena-Lite outperforms the baseline-mediated method.
+Note that Arena-Lite shows similar or superior reliability even in extreme data-poor benchmark condition ($|X|=50$).
 Table 3 presents the outcomes when using other LLMs as judges, with a fixed number of prompts ($|X|=500$).
 The results for Claude3.5-sonnet, Llama3.1-8b, and Qwen2.5-7b follow a similar trend.
-However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable for benchmarking.
-Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using VARCO Arena or considering reference-based approach.
+However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable as an LLM judge.
+Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using Arena-Lite or considering baseline-mediated approach.
 Spearman corr. ($\uparrow$) $|X|$ = 50 100 250 475 500
-comp. to ref. (4o) 0.895 0.935 0.963 0.966 0.964
-tournament (4o) 0.905 0.940 0.960 0.970 0.970
-comp. to ref. (4o-mini) 0.895 0.908 0.917 0.916 0.912
-tournament (4o-mini) 0.901 0.919 0.931 0.933 0.933
-Table 2: Spearman correlation ($\uparrow$) varying over size of the benchmark set ($|X|$) for each benchmarking approach. Comp. to ref. refers to reference-based approach.
-$|X|$ = 500 claude3.5 sonnet llama3.1 8b-it qwen2.5 7b-it qwen2.5 0.5b-it gemma2 2b-it
-comp. to ref. 0.924 0.820 0.756 0.089 0.592
-tournament 0.930 0.850 0.811 -0.124 0.552
-Table 3: Spearman correlation ($\uparrow$) result using other LLMs as a judge. Comp. to ref. refers to reference-based approach.
+baseline-mediated (4o) 0.895 0.935 0.963 0.966 0.964
+Arena-Lite (4o) 0.905 0.940 0.960 0.970 0.970
+baseline-mediated (4o-mini) 0.895 0.908 0.917 0.916 0.912
+Arena-Lite (4o-mini) 0.901 0.919 0.931 0.933 0.933
+Table 2: Robustness of ranking methods to benchmark set size, $|X|$ (Experiment 2, Sec.
+4.3).
+The table shows the median Spearman correlation ($\uparrow$) from 500 trials.
+Arena-Lite consistently achieves higher correlation than the baseline-mediated approach across all dataset sizes ($|X|$), demonstrating its superior reliability and robustness for ranking LLMs.
+$|X|$ = 500
+claude3.5
+sonnet
+llama3.1
+8b-it
+qwen2.5
+7b-it
+qwen2.5
+0.5b-it
+gemma2
+2b-it
+baseline-mediated 0.924 0.820 0.756 0.089 0.592
+Arena-Lite 0.930 0.850 0.811 -0.124 0.552
+Table 3: Robustness of ranking methods to the choice of judge LLM (Experiment 2, Sec.
+4.3).
+The table shows the Spearman correlation ($\uparrow$) between ground-truth LLM rankings and the results from each method.
+The values are medians from 500 trials.
+The results suggest that around 7B parameters-large LLMs is a viable minimum threshold for a reliable judge.
+Full results for other benchmark sizes are available in Appendix, Table 4.
 5.3 Incorporating a New LLM into an Existing Leaderboard
-While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case for leaderboards.
-We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a reference.
-Our findings indicate that the latter approach is more reliable (Table 4).
+While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case.
+We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a baseline.
+Our findings indicate that the later approach is more reliable (Table 5, Appendix).
 Further details and discussions are provided in Appendix A.6.
-$\bar{|\Delta_{\text{rank}}|}$ ($\downarrow$) gt=1-6 7-13 14-19 (20) total avg.
-binary search (4o) 0.92 1.84 2.13 1.72
-comp. to 1st (4o) 1.98 1.55 1.57 1.39
-binary search (4o-mini) 1.27 1.82 1.21 1.5
-comp. to 1st (4o-mini) 1.00 1.43 1.43 1.37
-Table 4: Comparison of the binary search method versus using the top-performing model’s response as a reference (comp.
-to 1st) for inserting a new LLM into the leaderboard.
-We report the mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from the ground-truth leaderboard as an additional error metric.
-For further details, see Algorithm 2 in Appendix.
 6 Related Works
-6.1 Elo-based LLM Benchmarking
-Recent studies have leveraged Elo ratings derived from human preferences to benchmark LLMs Boubdir et al.
-(2023).
-For instance, Chatbot Arena Chiang et al.
-(2024) employs Elo as an evaluation metric, while RAGElo Rackauckas et al.
-(2024) uses it as a ranking metric.
-These implementations highlight the benefits of applying Elo ratings for open-ended text quality evaluation.
-6.2 Reference-free Evaluation
-Reference-free evaluation has emerged as a promising alternative to static reference-based methods from the era of Natural Language Generation.
-Advances in LLM capabilities have enabled models to assess open-ended responses effectively Jauhiainen and Guerra (2024).
-When reference quality is poor, reference-free metrics like XComet Guerreiro et al.
-(2023) works as a more reliable alternatives.
+6.1 LLM-as-a-Judge for Systems Ranking
+Utilizing LLM-as-a-Judge as a building block for systems ranking has become a common practice in the LLM benchmarking community.
+Several studies have investigated how LLM judges compare to human evaluators, examining their similarities and differences Park et al.
+(2024), as well as how these differences impact system rankings (e.g., JuStRank Gera et al.
+(2024)Gao et al.
+(2025)).
+Our research extends these approaches by proposing a method that orchestrates LLM-as-a-Judge through a well-established tournament structure to derive rankings among systems.
+6.2 Efficient and Reliable Evaluation
+There is a growing body of research focused on optimizing the number of evaluations while maintaining reliability when using LLM-as-a-Judge for system ranking.
+Perlitz et al.
+proposed a metric called DIoR to quantify the relationship between computational costs and system ranking reliability.
+UniCBE Yuan et al.
+(2025) introduced a method to analyze the relationship between reliability and the number of judge evaluations based on uncertainty.
+BenchBench Perlitz et al.
+(2024b) systematically analyzed consistency across benchmarks and provided a package to facilitate this analysis.
+tinyBenchmarks Maia Polo et al.
+(2024) explored strategies to minimize the number of evaluations across various established benchmarks.
+Arena-Lite relates to these studies in that it leverages the properties of tournament structures and direct comparisons to achieve more reliable results with fewer judge evaluations.
 7 Conclusion
-We introduced VARCO Arena, a reference-free LLM benchmarking approach that employs a tournament-style framework with direct pairwise comparisons to evaluate the generative capabilities of LLMs.
-VARCO Arena offers a cost-effective, scalable, and adaptable solution for benchmarking response quality.
-Our simulation study and empirical experiments demonstrate that VARCO Arena consistently achieves higher rank reliability compared to current reference-based methods, as evidenced by stronger correlations with human-established Elo ratings.
-Given its robust performance and flexibility, we believe VARCO Arena can serve as a reliable automated tool for model selection and ranking across diverse and evolving use cases.
-Future work will explore broader applications, such as benchmarking LLMs in multi-modal settings (e.g., those incorporating visual or audio inputs and outputs).
+We introduced Arena-Lite, an efficient and reliable framework for evaluating Large Language Models (LLMs) through tournament-based direct comparisons.
+By eliminating the need for baseline encore outputs and adopting head-to-head comparison, Arena-Lite achieves higher reliability in system rankings with reduced number of comparisons.
+Our experiments, encompassing controlled stochastic modeling and empirical validation with various LLM judges, confirm that Arena-Lite consistently outperforms standard baseline-mediated evaluation methods, even with smaller datasets or weaker judges.
+The release of an accessible web demonstration and code supports the adoption of Arena-Lite to help streamlining model development cycle across research and industry.
+Future work will extend Arena-Lite’s application to diverse domains, including multi-modal LLM evaluation involving visual or audio inputs and outputs.
 Limitations
-Although we tested robustness of the tournaments performed by VARCO Arena, it still has added factor for randomness such as match bracket initialization which does not applies to reference-based method.
-Rooms for improvement exist for more informative match-making algorithm that would achieve better ranking than single-elimination tournaments within same or less number of matches.
+While we conducted extensive testing to assess the robustness of Arena-Lite tournaments—including 50 and 500 trials for Experiment 1 and Experiment 2, respectively—some inherent sources of randomness remain, such as variation due to initial match bracket assignments.
+The randomness in bracket assignment is added for adopting tournament structure of Arena-Lite and may influence outcome stability.
+Future work could explore more informative or adaptive matchmaking strategies that improve ranking fidelity beyond what is achievable with single-elimination formats, potentially within the same or even fewer number of matches.
 Appendix A Appendix
-A.1 Machine Requirements for Experiments
+A.1 Arena-Lite Web Demo
+We provide screenshots of Arena-Lite web demo here (link to Huggingface Space).
+You could try or locally host Arena-Lite according to its documentation.
+It is quite easy to use and do not require much of resources to host.
+Arena-Lite provides the benchmark result (Figure 4) with helpful visualization interface that enables walking through the matches and tournaments one by one (Figure 5) and match statistics between LLMs (Figure 6).
+We also provide visualization that helps examining potential bias of LLM Judge being used (Figure 7).
+Figure 4: Arena-Lite web screenshot 1: At the top of the result page, one can see the leaderboard of LLMs with their BT preference.
+If the benchmark dataset has subcategories, radar chart (right) is also visible.
+Figure 5: Arena-Lite web screenshot 2: User can walk through the matches and tournaments one by one.
+Match brackets is visualized briefly with text UI and user can select any specific match to see the details (e.g.
+match result, prompt, and model outputs).
+Figure 6: Arena-Lite web screenshot 3: User can see the match statistics between LLMs (i.e. win rate between model pairs, number of matches per pair and per model).
+Figure 7: Arena-Lite web screenshot 4: User can see the LLM Judge’s examine how biased the LLM judge being used. The demo provides clues for potential bias toward response length and position.
+A.1.1 Starter Prompt Set
+We provide several judge prompts that we have used for specific target tasks.
+Some of those are for quite specialized tasks, and some might work for evaluating general instruction following.
+You could customize your own judge prompt based on your evaluation needs according to the documentation.
+The list of the judge prompts we provide is as follows, and one could see the detailed prompts at yaml files here.
+  1. 1.
+    llmbar prompt (Figure A.8.2) and conciser version of the prompt, llmbar_brief. Those are for evaluating instruction following.
+  2. 2.
+    translation_pair prompt for selecting translation models trained on game-specialized parallel corpora,
+  3. 3.
+    rag_pair_kr prompt for evaluating knowledge groundedness of korean RAG models over chatting scenario,
+  4. 4.
+    translation_fortunecookie prompt which was crafted for evaluating translation models specialized for translating fortune-tellings, and
+  5. 5.
+    post_edit prompt for evaluating conversation revision based on given persona of a speakers given in a instruction, which is quite specialized use case.
+A.2 Full table for Experiment 2
+Here is the extended results of Experiment 2 (Section 4.3) presented in Table 3.
+Aligned LLMs smaller than 7B parameters struggles to work as a proper Judge.
+Otherwise, Arena-Lite method excels over common practice of using encore outputs as baselines.
+Table 4: Extended results for comparing Arena-Lite to baseline method of using encore outputs. We tested other LLMs as judge over various size of benchmark datasets.
+Dataset size method claude 3.5 sonnet llama3.1-8b-it qwen2.5-7b-it qwen2.5-0.5b-it gemma2-2b-it
+50 baseline-mediated .896 .656 .492 .010 .064
+Arena-Lite (ours) .897 .715 .544 -0.051 -0.088
+100 baseline-mediated .912 .732 .596 .002 .079
+Arena-Lite (ours) .918 .780 .656 -0.068 -0.090
+250 baseline-mediated .924 .801 .700 .045 .560
+Arena-Lite (ours) .929 .830 .760 -0.131 .551
+475 baseline-mediated .924 .819 .708 .083 .112
+Arena-Lite (ours) .930 .845 .810 -0.131 -0.009
+500 baseline-mediated .924 .820 .756 .089 .592
+Arena-Lite (ours) .930 .850 .811 -0.124 .551
+A.3 Machine Requirements for Experiments
 Except the part we inferenced open-weight models such as Llama, Qwen and Gemma, our experiments are mostly do not require GPU usage.
@@
 The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
-A.2 Assuring Statistical Significance of the Results within Budget for proprietary models
+A.4 Assuring Statistical Significance of the Results within Budget for proprietary models
 To ensure a statistically significant number of trials for each experiment while staying within budget, we utilize OpenAI’s Batch API to prepare full-grid match outcomes (i.e., all-play-all matches for every prompt) in a cache file, allowing us to reuse these outcomes.
@@
 When experimenting with a subset of the Arena-Hard-Auto benchmark ($|X|<500$), we sample a stratified subset of the benchmark dataset for each new trial.
-Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al.
+Figure 8: BT preference of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al.
 (2024) prompts.
-VARCO Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category).
+Arena-Lite result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category).
 Error bars are 95% confidence intervals.
-A.3 Elo ratings from VARCO Arena compared to Human Annotations
-Figure 4 shows the Elo ratings computed out of VARCO Arena.
+A.5 BT preference from Arena-Lite compared to Human Annotations
+Figure 8 shows the BT preference computed out of Arena-Lite.
 For judge, we used gpt-4o.
-As mentioned in the caption, the Elo ratings are bootstrapped median value from 500 trials.
+As mentioned in the caption, the BT preference are bootstrapped median value from 500 trials.
 95% confidence intervals also plotted as an error bar, which look negligible in scale compared to observed values.
 Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
-A.4 Binary search vs. Win rate over reference
+A.6 Binary search vs. Win rate over baseline
 Algorithm 2 Binary Search for Enlisting new LLM to a leaderboard
@@
 2:Updated leaderboard $L^{\prime}$ with $m_{\text{new}}$ placed
-3:$n_{\text{comparisons}}\leftarrow\lfloor\log_{2}(|L|)\rfloor$
-4:$n_{\text{matches}}\leftarrow\lfloor|X|/n_{\text{comparisons}}\rfloor$
-5:function BinarySearchPlacement($L,m_{\text{new}}$)
+[…]
+[…]
+[…]
 6:  X $\leftarrow$ Shuffle(X)
 7:  X $\leftarrow$ concat(X;X)
-8:     low $\leftarrow 0$
+8:  low 0$\leftarrow$ 0← 0
 9:  high $\leftarrow|L|-1$
 10:  while low $\leq$ high do
-11:         mid $\leftarrow\lfloor(\text{low}+\text{high})/2\rfloor$
-12:         wins $\leftarrow 0$
+[…]
+12:   wins 0$\leftarrow$ 0← 0
 13:   for $i\leftarrow 1$ to $n_{\text{matches}}$ do
@@
 […]
-16:                  wins $\leftarrow$ wins $+1$
+16:      wins $\leftarrow$ wins +1+1+ 1
 17:     end if
@@
 19:   if wins $>n_{\text{matches}}/2$ then
-20:              high $\leftarrow$ mid $-1$
+20:     high $\leftarrow$ mid 1-1- 1
 21:   else if wins $<n_{\text{matches}}/2$ then
-22:              low $\leftarrow$ mid $+1$
+22:     low $\leftarrow$ mid +1+1+ 1
 23:   else if $|X|>$then
@@
 30:end function
-31:function UpdateLeaderboard($L,m_{\text{new}}$)
-32:     position, istie $\leftarrow$ BinarySearchPlacement($L,m_{\text{new}}$)
+[…]
+[…]
 33:  $L^{\prime}\leftarrow L$.insert(position, $m_{\text{new}}$, istie)
@@
 35:end function
-A.4.1 Binary Search
-We tried binary search placement of a newly added LLM to the leaderboard without reference text in Table 5.
-Details of how we implemented binary search are attached in Appendix 2.
-It turns out that binary search based on leaderboard ranks is not as reliable as the current approach of scoring the newcomer to the reference outputs.
-The number of judge operations performed is equivalent to the matches allocated to the least-performant model in a tournament, which is $|X|$ (i.e.
-maximum possible matches that an LLM could have is $|X|*\text{log}_{2}n_{\text{model}}$).
-Within the size of the benchmark prompts ($|X|$), binary search is incompatible with the current approach of using reference instead.
-A.4.2 Comparing to the most performant Model so far: Converting Elo Table back to Win Rate
+A.6.1 Binary Search
+We tried binary search placement of a newly added LLM to the leaderboard without baseline output in Table 6.
+Details of how we implemented binary search are attached in Algorithm 2, Appendix.
+It turns out that binary search based on already built leaderboard ranks is not as reliable compared to utilizing the best model’s outputs as a baseline.
+Therefore, when adding a newcomer LLM to pre-existent leaderboard, we could utilize the already submitted responses as a baseline from the 1st placed LLM.
+$\bar{|\Delta_{\text{rank}}|}$ ($\downarrow$) gt=1-6 7-13 14-19 (20) total avg.
+binary search (4o) 0.92 1.84 2.13 1.72
+comp. to 1st (4o) 1.98 1.55 1.57 1.39
+binary search (4o-mini) 1.27 1.82 1.21 1.5
+comp. to 1st (4o-mini) 1.00 1.43 1.43 1.37
+Table 5: Comparison of the binary search method versus using the top-performing model’s response as a baseline (comp.
+to 1st) for inserting a new LLM into the leaderboard.
+We report the mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from the ground-truth leaderboard as an additional error metric.
+For further details, see Algorithm 2 in Appendix.
+A.6.2 Comparing to the most Performant Model so far: Converting Ratings Table back to Win Rates
 Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion.
-One could pick an appropriate anchor LLM as a reference in a leaderboard to estimate the skill of a newcomer.
-Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the reference.
-This is the same as converting the Elo table into a win rate leaderboard.
-Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the reference regardless we have picked.
-There are two ways to estimate the win rate over the reference model.
-We could just count the matches given are enough in amount, or we could also convert Elo ratings back to $P(i>a)$ to use it directly for scoring for the model ranks in the leaderboard.
-Reminding that Elo rating is purposed for expecting a likely outcome of the match, this should work.
+One could pick an appropriate anchor LLM as a baseline in a leaderboard to estimate the skill of a newcomer.
+Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the baseline.
+This is the same as converting the preference ratings table into a win rate leaderboard.
+Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the baseline regardless we have picked.
+There are two ways to estimate the win rate over the baseline model.
+We could just count the matches given are enough in amount, or we could also convert BT preference back to $P(i>a)$ to use it directly for scoring for the model ranks in the leaderboard.
+Reminding that BT preference rating is for expecting a likely outcome of the match, this should work.
 After this win rate of the newcomer model $P^{*}(n>a)=\frac{\text{count(n wins)}}{|X|}$ could be directly compared for enlisting.
@@
 (2024)).
-A.5 Separability In terms of Confidence Interval
+A.7 Separability In terms of Confidence Interval
 To see how well the two benchmarking approach (anchored comparison and tournament approach) separates LLMs in adjacent ranks, we provide scatter plot of Elo rating and win rate paired with error bar (95% confidence interval).
@@
 Figure 10: gpt-4o result of anchored comparison and tournament approach. 1000 bootstrapped median from 500 observations used for confidence interval estimation.
-A.6 Judge configuration
-A.6.1 Evaluation Prompt
+A.8 Judge configuration
+A.8.1 Evaluation Prompt
 We use the prompt from LLMBar.
@@
 You can refer to the original prompt in LLMBar github.
-A.6.2 Decoding Parameters
+A.8.2 Decoding Parameters
 We did not configure decoding parameters of judge LLMs (gpt-4o[-mini]), which its temperature defaults to 1.